Krell Industries has a share price of $ 22.23 today. If Krell is expected to pay a dividend of $ 1.11 this year and its stock price is expected to grow to $ 23.04 at the end of the​ year, what is​ Krell's dividend yield and equity cost of​ capital? The dividend yield is nothing​%. ​(Round to one decimal​ place.)
Answer #1

Dividend yield=Expected dividend/Beginning value

=(1.11/22.23)=5%(Approx).

Equity cost of capital=(End value-Beginning value+Dividend)/Beginning value

=(23.04-22.23+1.11)/22.23

=8.6%(Approx).